Week of 201116

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, DB), Julia (WLCG), Maarten (WLCG, ALICE), Eric (storage), Ivan (ATLAS), Darren (RAL), DaveM (FNAL), Luis FA (computing), Andrew (TRIUMF), Borja (monitoring), Xin (BNL), MarkS (LHCb), Priyanshu (computing), Christoph (CMS), Elena (CNAF), DavidB (IN2P3), Pepe (PIC), Maurizio (DB)

Experiments round table:

  • CMS reports ( raw view) -
    • Again good CPU utilization (~200k cores for production & ~50k cores for analysis)
    • CMS re-establishes Loadtests with Rucio
      • A couple of tickets were opened to request write access to LFN space /store/test/loadtest

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • IN2P3 dCache namespace migration to be performed 17/11/20 (tomorrow) without downtime
    • Activity:
      • MC and WG productions; User jobs.

Sites / Services round table:

  • ASGC: nc
  • BNL: T1 downtime on Dec 1st, 6am~6pm EST, due to site-wide network intervention/upgrade.
  • CNAF: NTR
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR:NTR
  • KISTI: nc
  • KIT: A downtime is declared for tomorrow 10 a.m. ( GOCDB:29747), where we will restart all LHCb dCache pools. This is necessary for enabling TPC over xrootd for them.
  • NDGF: nc
  • NL-T1:
  • NRC-KI: nc
  • OSG: nc
  • PIC: There was downtime over the weekend due to electrical maintenance. All services should be fully recovered now.
  • RAL: NTR
  • TRIUMF: dCache downtime, Wednesday November 18 UTC 18:00hrs - 23:00hrs: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=29722

  • CERN computing services:
    • The CE upgrades to CentOS7 are ongoing. To be finished next week.
    • Network interventions may impact batch jobs. Today: OTG0060215 Wednesday: OTG0060216.
    • Loss of capacity in Point8 this morning due to power incident: OTG0060381.
  • CERN storage services:
  • CERN databases: PDBR database will be updated to Oracle 19c next Monday. Downtime of up to 3h (OTG:0060397)
  • GGUS: NTR
  • Monitoring: NTR
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2020-11-16 - KateDziedziniewicz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback